OpenStudy (anonymous):

Susan is creating a solution. A formula for the solution calls for 7/10 mL of water. Susan wants to make 2/7 less of the solution than the formula creates. How much water should Susan use? Express your answer as a fraction in simplest form? Can someone please explain to me why the solution is 1/2?

OpenStudy (stamp):

Susan wants 2/7 less of the solution, which means she wants 5/7 total. Take 5/7 of 7/10 mL \[\frac{5}{7}*\frac{7}{10}=\frac{5}{10}=\frac{1}{2}\]

OpenStudy (anonymous):

so basically 7/7 - 2/7 would be 5/7?